MIDC/Data Quality Statements
Total Records: 9, displaying 1-9

R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSDirect PGS-10005/30/201416:18:0005/31/201411:30:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
2BMSDirect PGS-10005/01/201313:27:0005/02/201306:40:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
3BMSDirect PGS-10007/06/201219:35:0007/09/201210:18:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
4BMSDirect PGS-10007/12/201122:57:0007/13/201107:40:00Due to a tripped GFCI outlet at SRRL the following instruments were without power during this time: TWC & POM-01 Photometers, Direct CUVA2 & CUVB2, and the PGS100. The readings are either zero or bogus and should not be used. GFCI was tripped due to Eppley venitlators which were water logged by thunderstorm.
5BMSS09032Direct PGS-10007/04/201018:10:0007/06/201007:30:00Due to several tripped GFCI outlets at SRRL the following instruments were without power from July 4 @ 18:10 to July 6 @ 7:30 MST: Global CUVA1, CUVB1, MS210W, 501A and Direct CUVA2, CUVB2, PGS100. The readings are either zero or bogus and should not be used.
6BMSS08029Direct PGS-10004/27/201008:00:0005/06/201012:00:00The NREL owned PGS-100 was returned to service April 22 at 10:50 MST. Factory found no problem with the shutter. However, the shutter is sticking closed again (causing no signal) on April 27. The loaner instrument was installed again at 12:20 MST on 5/6/2010. The loaner (S/N S09032) is now owned by NREL and the old unit (S/N S08029) will be returned to manufacturer.
7BMSS09032Direct PGS-10001/29/201014:40:0004/22/201010:45:00While the NREL PGS-100 unit is out for repair, the factory supplied NREL with a loaner unit. This unit will verify the operation of the data acquisition, Prede Solar Tracker, cabling, etc. Data is primarily for functional check. Instrument was calibrated by NREL on Feb 26, 2010 outdoors against the NIST traceable LI-1800 PRS-158. New calibration file was applied on March 11, 2010 and all historical data for the loaner unit was corrected. The loaner sensor and control box worked with no problems.
8BMSS08029Direct PGS-10011/18/200900:00:0001/29/201014:35:00The SRRL PGS-100 Direct Normal Spectroradiometer shutter is sticking closed again (causing no signal). First intermittent occurrence was on 11/18/09 and problem has been getting worse as time goes on. Instrument was removed from service on 1/21/2010 and will be sent back to factory again.
9BMSS08029Direct PGS-10007/08/200900:00:0011/10/200913:49:00The shutter mechanism started having problems on July 8, 2009, the result is that many spectra would read zero for all wavelengths. This was an intermittent problem and later it started failing more frequently. It was removed from service on September 28, 2009 and sent back to the factory in Japan for repair. Instrument was returned to service on November 10, 2009 at 13:50 MST. The shutter solenoid was replaced at the factory.